a spaceship is designed with a rotating section to stimulate the effect of gravity for the astronauts. The internal radius of this section is 18.0 m. Calculate the tangential speed that the other wall must have in order for the astronauts to experience 1/2 g of artificial gravity when standing on it?


1
Expert's answer
2022-03-07T12:21:54-0500

The centripetal acceleration

"a=\frac{v^2}{R}=\frac{1}{2}g"

The tangential speed

"v=\sqrt{\frac{R}{2}g}=\sqrt{\frac{18.0}{2}*9.8}=9.4\:\rm m/s"